Determine the equation of combined variation. Then solve for the missing value.
x varies directly with y and inversely with z.
x = 5 when y = 10 and z = 5.
Find x when y = 20 and z = 10.

Answer by stanbon(75887)

Determine the equation of combined variation. Then solve for the missing value.
------------------------------------------
x varies directly with y and inversely with z.
x = k*y/z
Solve for "k" if x = 5 when y = 10 and z = 5.
5 = k*10/5
k = 2.5
----
Equation:
x = 2.5*y/z
----
Find x when y = 20 and z = 10.
x = 2.5*20/10
x = 5
---------------
Cheers,
Stan H.
----------------